National Defence Academy (NDA) · Topic Collection

Quantitative Aptitude PYQs

Arithmetic, number systems, percentages, ratios, profit and loss, time and work, algebra, geometry, mensuration, and trigonometry.

2520 Total Questions21 Official PapersYears: 2015–2025Search in Quantitative Aptitude
Showing questions 581–600 of 2520Page 30 of 100
NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 581

National Defence Academy (NDA) 2023• Mathematics• Mathematics

The sum of deviations of n numbers from 10 and 20 are p and q respectively. If (p - q)2 = 10000, then what is the value of n?

  1. AOption (a)

    10

  2. BOption (b)

    20

  3. COption (c)

    50

  4. DOption (d)

    100

View answer & explanation

Correct answer

Answer: (a)A.

10

Why this answer is correct

Given: 

The sum of deviations of n numbers from 10 = p

The sum of deviations of n numbers from 20 = q

(p - q)2 = 10000

Calculation:

p = \(\displaystyle \sum {(x_i-10)} = \displaystyle \sum {x_i-10n}\)

q = \(\displaystyle \sum {(x_i-20)} = \displaystyle \sum {x_i-20n}\)

As, (p - q)2 = 10000

\((\displaystyle \sum {x_i-10n} - \displaystyle \sum {x_i+20n})^2 = 10000\)

\((10n)^2 = 10000\)

n = 10

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 582

National Defence Academy (NDA) 2023• Mathematics• Mathematics

If X̅ = 20 is the mean of 10 observations x1, x2, ... x10; then what is the value of \(\displaystyle \sum_{i=1}^{10}\left(\frac{3 x_i-4}{5}\right) ?\) ?

  1. AOption (a)

    0

  2. BOption (b)

    12

  3. COption (c)

    112

  4. DOption (d)

    1012

View answer & explanation

Correct answer

Answer: (c)C.

112

Why this answer is correct

Given:

Mean (X̅) = 20, No. of observations = 10

Calculation:

Let

S = \(\displaystyle \sum_{i=1}^{10}\left(\frac{3 x_i-4}{5}\right)\)

S = \(\displaystyle \sum_{i=1}^{10}\left(\frac{3 x_i}{5}\right) -{4\over 5} \times \displaystyle \sum_{i=1}^{10} 1\)

S = \({3\over 5} \displaystyle \sum_{i=1}^{10}( x_i) -{4\over 5} \times 10\)           -----(1)

As, \({{\displaystyle \sum_{i=1}^{10}( x_i)}\over 10} = 20\)

\({\displaystyle \sum_{i=1}^{10}( x_i)} = 200\)

Now, from equation (1)

S = \({3\over 5} (200) -{4\over 5}(10)\)

S = \({600-40\over 5}\)

S = 112

∴ \(\displaystyle \sum_{i=1}^{10}\left(\frac{3 x_i-4}{5}\right)\) = 112

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 583

National Defence Academy (NDA) 2023• Mathematics• Mathematics

If the mean and the sum of squares of 10 observations are 40 and 16160 respectively, then what is the standard deviation?

  1. AOption (a)

    16

  2. BOption (b)

    6

  3. COption (c)

    5

  4. DOption (d)

    4

View answer & explanation

Correct answer

Answer: (d)D.

4

Why this answer is correct

Formula Used:

Standard Deviation = \(\sigma = \sqrt{({\sum x_i\over n})^2 -({\sum x^2_i\over n}) }\)

Given:

Mean = 40, Sum of squares = 16160, No. of observations (n) = 10

\({\sum x_i\over n} = 40\), \(\sum x^2_i = 16160\)

Calculation:

\(\sigma = \sqrt{({\sum x^2_i\over n}) -({\sum x_i\over n})^2}\)

\(\sigma = \sqrt{{16160\over 10} - 40^2}\)

\(\sigma = \sqrt{{1616} - 1600}\)

\(\sigma = \sqrt{16}\)

\(\sigma = 4\)

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 584

National Defence Academy (NDA) 2023• Mathematics• Mathematics

Three dice are thrown. What is the probability of getting a sum which is a perfect square?

  1. AOption (a)

    \(\frac{17}{108}\)

  2. BOption (b)

    \(\frac{5}{108}\)

  3. COption (c)

    \(\frac{19}{108} \)

  4. DOption (d)

    \(\frac{23}{108}\)

View answer & explanation

Correct answer

Answer: (a)A.

\(\frac{17}{108}\)

Why this answer is correct

Concept:

Perfect Squares = 1, 4, 9, 16, 25, 36

Among all these, we can get perfect squares 4, 9, 16

Calculation:

The total number of outcomes = 63 = 216

Total ways for an outcome of 4,

(1, 1, 2) in 3 ways

Total ways for an outcome of 9,

(1, 2, 6) in 3! ways

(1, 3, 5) in 3! ways

(2, 3, 4) in 3! ways

(3, 3, 3) in 1! ways

(2, 2, 5) in 3 ways

(1, 4, 4) in 3 ways

Total Ways =  3! + 3! + 3! + 1! + 3 + 3 = 6 + 6 + 6 + 1 + 3 + 3 = 25

Total ways for an outcome of 16,

(4, 6, 6) in 3 ways

(5, 5, 6) in 3 ways

Total ways = 3 + 3 = 6

Number of favorable outcomes = 3 + 25 + 6 = 34

The probability of getting a sum which is a perfect square,

= 34/216

= 17/108

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 585

National Defence Academy (NDA) 2023• Mathematics• Mathematics

A, B, C and D are mutually exclusive and exhaustive events.

If 2P(A) = 3P(B) = 4P(C) = 5P(D), then what is 77P(A) equal to ?

  1. AOption (a)

    12

  2. BOption (b)

    15

  3. COption (c)

    20

  4. DOption (d)

    30

View answer & explanation

Correct answer

Answer: (d)D.

30

Why this answer is correct

Concept: 

Mutually exclusive events are those events that do not occur at the same time.

For example, when a coin is tossed then the result will be either head or tail, but we cannot get both the results. 

Sum of mutually exclusive events is equals to 1.

Calculation:

A, B, C and D are mutually exclusive and exhaustive events. Therefore,

P(A) + P(B) + P(C) + P(D) = 1         -----(1)

According to the question,

2P(A) = 3P(B) = 4P(C) = 5P(D) = k (Let)

P(A) = k/2, P(B) = k/3, P(C) = k/4, P(D) = k/5

Put the values in equation (1),

P(A) + P(B) + P(C) + P(D) = 1

\({k\over 2} + {k\over 3} + {k\over 4} + {k\over 5} = 1\)

\({30k + 20k+15k+12k\over 60}= 1\)

\({77k}= 60\)

Put k = 2 P(A),

77 P(A) = 30

Hence, 77P(A) equals to 30.

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 586

National Defence Academy (NDA) 2023• Mathematics• Mathematics

Two distinct natural numbers from 1 to 9 are picked at random. What is the probability that their product has 1 in its unit place?

  1. AOption (a)

    \(\frac{1}{81}\)

  2. BOption (b)

    \(\frac{1}{72}\)

  3. COption (c)

    \(\frac{1}{18}\)

  4. DOption (d)

    \(\frac{1}{36}\)

View answer & explanation

Correct answer

Answer: (d)D.

\(\frac{1}{36}\)

Why this answer is correct

Given:

Two distinct natural numbers from 1 to 9 are picked at random.

Calculation:

Total number of ways in which two distinct natural numbers from 1 to 9 are picked at random = 9C2

9C2 = \({{9!}\over {{(9-2)!}2!}} \) = 36

Favourable pairs are 

 (7, 3)

Total favourable pairs = 1

Required probability = 1/36

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 587

National Defence Academy (NDA) 2023• Mathematics• Mathematics

Two dice are thrown. What is the probability that difference of numbers on them is 2 or 3 ?

  1. AOption (a)

    \(\frac{7}{36}\)

  2. BOption (b)

    \(\frac{7}{18}\)

  3. COption (c)

    \(\frac{5}{18}\)

  4. DOption (d)

    \(\frac{11}{36}\)

View answer & explanation

Correct answer

Answer: (b)B.

\(\frac{7}{18}\)

Why this answer is correct

Calculation:

Sample Space:

(1,1); (1,2); (1,3); (1,4); (1,5); (1,6);

(2,1); (2,2); (2,3); (2,4); (2,5); (2,6);

(3,1); (3,2); (3,3); (3,4); (3,5); (3,6);

(4,1); (4,2); (4,3); (4,4); (4,5); (4,6);

(5,1); (5,2); (5,3); (5,4); (5,5); (5,6);

(6,1); (6,2); (6,3); (6,4); (6,5); (6,6);

From the above sample space, there are 14 possibilities for the difference of numbers between them is 2 or 3

Total number of Sample Space = 36

the probability that the difference of numbers on them is 2 or 3,

= 14/36

= 7/18

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 588

National Defence Academy (NDA) 2023• Mathematics• Mathematics

What is the mean of the numbers 1, 2, 3, ... 10 with frequencies 9C0, 9C1, 9C2 ..., 9C9, respectively?

  1. AOption (a)

    1.1 × 28

  2. BOption (b)

    1.2 × 74

  3. COption (c)

    5.5

  4. DOption (d)

    0.55

View answer & explanation

Correct answer

Answer: (c)C.

5.5

Why this answer is correct

Given:

Frequencies (fi) = 9C0, 9C1, 9C2 ..., 9C9

Numbers (xi) = 1, 2, 3, ... 10

Formula Used:

\(Mean = {{\sum f_ix_i}\over {\sum f_i}}\)

\(:^nC_0+: 2\times^nC_1 + 3\times:^nC_2 + 4\times:^nC_3 .......+^nC_n (n) = (n+2)(2)^{n-1}\)

\(:^nC_0+:^nC_1 + :^nC_2 .......+^nC_n = (2)^n\)

Calculation:

\(Mean = {{1 \times^9C_0 + 2 \times^9C_1+3 \times^9C_2.....10 \times^9C_9}\over {^9C_0 + ^9C_1+^9C_2.....^9C_9}}\)

\(Mean = {{(9+2)\times {2^{9-1}}}\over {2^9}}\)

\(Mean = {{(11)\times {2^{8}}}\over {2^9}}\)

Mean = 5.5

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 589

National Defence Academy (NDA) 2023• Mathematics• Mathematics

The probability that a person recovers from a disease is 0.8. What is the probability that exactly 2 persons out of 5 will recover from the disease?

  1. AOption (a)

    0.00512

  2. BOption (b)

    0.02048

  3. COption (c)

    0.2048

  4. DOption (d)

    0.0512

View answer & explanation

Correct answer

Answer: (d)D.

0.0512

Why this answer is correct

Concept: 

Binomial distribution is the discrete probability distribution that gives only two possible results in an experiment, either Success or Failure. For example, if we toss a coin, there could be only two possible outcomes: heads or tails, and if any test is taken, then there could be only two results: pass or fail. This distribution is also called a binomial probability distribution.

P (x) = \(^{n}C_x \times [P(Success)]^x \times [P(Failure)]^{n-x}\)

Given:

Success: The probability that a person recovers from a disease is 0.8. 

Failure: The probability that a person does not recover from a disease = 1 - 0.8 = 0.2 

Calculation:

P (Success) = 0.8, P(Failure) = 0.2, n = 5, x = 2

P (x = 2) = \(^{5}C_2 \times (0.8)^2 \times(0.2)^{5-2}\)

P (x = 2) = \({{5!}\over {{(5-2)!}2!}} \times (0.64) \times(0.008)\)

P (x = 2) = \(10 \times (0.00512)\)

P (x = 2) = \(0.0512\)

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 590

National Defence Academy (NDA) 2023• Mathematics• Mathematics

Suppose that there is a chance for a newly constructed building to collapse, whether the design is faulty or not. The chance that the design is faulty is 10%. The chance that the building collapses is 95% if the design is faulty, otherwise it is 45%. If it is seen that the building has collapsed, then what is the probability that it is due to faulty design?

  1. AOption (a)

    0.10

  2. BOption (b)

    0.19

  3. COption (c)

    0.45

  4. DOption (d)

    0.95

View answer & explanation

Correct answer

Answer: (b)B.

0.19

Why this answer is correct

Concept:

Bayes theorem: It is used to determine the conditional probability of event A when event B has already happened. The general statement of Bayes’ theorem is “The conditional probability of an event A, given the occurrence of another event B, is equal to the product of the event of B, given A and the probability of A divided by the probability of event B.” i.e.

P(A|B) = \({{P(A) \times P(B|A)}\over {P(B)}}\)

Given:

Event A: The chance that the design is faulty.

Event B: The chance that the design is not faulty.

Event C: The chance that the building collapses

P(A) = 10%, P(B) = 90%, P(A|C) = 95%, P(B|C) = 45%

Calculation:

P(C|A) = \({{P(A) \times P(C|A)}\over {P(A) \times P(C|A)}+{P(B) \times P(C|B)}}\)

P(C|A) = \({{10 \times 95}\over {10 \times 95}+{90 \times 45}}\)

P(C|A) = \({{950}\over {950}+{4050}}\)

P(C|A) = \({{950}\over {5000}}\)

P(C|A) = 0.19

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 591

National Defence Academy (NDA) 2023• Mathematics• Mathematics

If r is the coefficient of correlation between x and y, then what is the correlation coefficient between (3x + 4) and (-3y + 3)?

  1. AOption (a)

    -r

  2. BOption (b)

    r

  3. COption (c)

    √3 r

  4. DOption (d)

    -√3r

View answer & explanation

Correct answer

Answer: (a)A.

-r

Why this answer is correct

Given: The correlation coefficient between x and y is r.

Concept:

The correlation coefficient measures the strength and direction of the linear relationship between two variables. 

Calculation:

The transformation of y involves multiplication by -3. Multiplying one of the variables by a negative value changes the direction of the correlation.

= - r 

Hence, the correlation coefficient between ( 3x + 4 ) and ( -3y + 3 ) is  - r

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 592

National Defence Academy (NDA) 2023• Mathematics• Mathematics

A fair coin is tossed 6 times. What is the probability of getting a result in the 6th toss which is different from those obtained in the first five tosses ?

  1. AOption (a)

    \(\frac{7}{16}\)

  2. BOption (b)

    \(\frac{1}{16}\)

  3. COption (c)

    \(\frac{1}{32}\)

  4. DOption (d)

    \(\frac{1}{64}\)

View answer & explanation

Correct answer

Answer: (c)C.

\(\frac{1}{32}\)

Why this answer is correct

Concept:

\(P(E) = {{n(E)} \over {n(S)}}\)

where n(E) = No. of favourable cases for event A and n(S) = cardinality of sample space.

If a coin is tossed then,

P(getting head) = \(1 \over 2\) P(getting tail) = \(1 \over 2\)

Calculation:

A fair coin is tossed repeatedly. Sixth toss is different from those obtained in the first five tosses.

There are two cases of getting the result as:

Case 1: Getting a tail is a result

Case 2: Getting a head is a result

The possibility is that in the first five tosses, we get 5 heads and 1 tail or five tails and one head,

⇒ \({1 \over 2} × {1 \over 2} × {1 \over 2} × {1 \over 2} × {1 \over 2} + {1 \over 2} ×{1 \over 2} ×{1 \over 2} ×{1 \over 2} ×{1 \over 2} \)

⇒ \({1 \over 64} + {1 \over 64}\)

⇒ \({2 \over 64}\) = \({1 \over 32}\)

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 593

National Defence Academy (NDA) 2023• Mathematics• Mathematics

If H is the Harmonic Mean of three numbers 10C4, 10C5, and 10C6, then what is the value of \(\frac{270}{H}\) ?

  1. AOption (a)

    1

  2. BOption (b)

    \(\frac{14}{17}\)

  3. COption (c)

    \(\frac{17}{14}\)

  4. DOption (d)

    \(\frac{1}{31}\)

View answer & explanation

Correct answer

Answer: (c)C.

\(\frac{17}{14}\)

Why this answer is correct

Given:

If H is the Harmonic Mean of three numbers 10C4, 10C5, and 10C6.

Concept: 

The harmonic mean is the reciprocal of the arithmetic mean of the values. It is calculated by dividing the number of observations by the reciprocal of each number in the series. 

\(^{n}C_k = {{n!}\over (n-k)! k!}\)

Calculation:

\(^{10}C_4 = {{10!}\over (10-4)! 4!} = 210\)

\(^{10}C_5 = {{10!}\over (10-5)! 5!} = 252\)

\(^{10}C_6 = {{10!}\over (10-6)! 6!} = 210\)

Harmonic mean (H) = \({{3} \over {{1\over210}+{1\over252}+{1\over210}}}\)

Harmonic mean (H) = \({{3} \over {{6+5+6\over1260}}}\)

Harmonic mean (H) = \({{3 \times 1260} \over {17}}\)

\(\frac{270}{H}\) = \({{17 \times270}\over {3 \times 1260}} = {17\over 14}\)

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 594

National Defence Academy (NDA) 2023• Mathematics• Mathematics

In a class, there are n students including the students P and Q. What is the probability that P and Q sit together if seats are assigned randomly?

  1. AOption (a)

    \(\frac{1}{n}\)

  2. BOption (b)

    \(\frac{2}{n}\)

  3. COption (c)

    \(\frac{4}{n}\)

  4. DOption (d)

    \(\frac{1}{2n}\)

View answer & explanation

Correct answer

Answer: (b)B.

\(\frac{2}{n}\)

Why this answer is correct

Given: In a class, there are n students including the students P and Q.

Calculation:

Total number of students in the class = n

The probability of P and Q sitting together:

Treat the P and Q as one unit (X).

Now, we have (n - 2, X) to arrange. This can be done in (n - 2 + 1)! ways.

Within the P and Q, they can arrange themselves in 2! ways.

Therefore, the total number of favorable outcomes in this case is (n - 1)! × 2!.

n people can be arranged in a row in n! ways.

\(Probability = {Favorable: outcomes \over Total :possible :outcomes}\)

Probability \(= {{((n - 1)! × 2!)} \over n!}\)

Probability \(= {{2} \over n}\)

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 595

National Defence Academy (NDA) 2023• Mathematics• Mathematics

In a Binomial distribution B(n, p), n = 6 and 9P(X = 4) = P(X = 2). What is p equal to ?

  1. AOption (a)

    \(\frac{1}{4}\)

  2. BOption (b)

    \(\frac{1}{2}\)

  3. COption (c)

    \(\frac{3}{4}\)

  4. DOption (d)

    \(\frac{4}{5}\)

View answer & explanation

Correct answer

Answer: (a)A.

\(\frac{1}{4}\)

Why this answer is correct

Concept: 

Binomial distribution is the discrete probability distribution that gives only two possible results in an experiment, either Success or Failure. For example, if we toss a coin, there could be only two possible outcomes: heads or tails, and if any test is taken, then there could be only two results: pass or fail. This distribution is also called a binomial probability distribution.

B(n, p) = \(^{n}C_x \times [P(Success)]^x \times [P(Failure)]^{n-x}\)

Given:

n = 6

9P(X = 4) = P(X = 2).

Calculation:

Let the success be p

Let the failure be 1 - p

P (x = 2) = \(^{6}C_2 \times (p)^2 \times(1-p)^{6-2}\)

P (x = 2) = \({{6!}\over {{(6-2)!}2!}} \times (p)^2 \times(1-p)^{4}\)

P (x = 2) = \(15 \times (p)^2 \times(1-p)^{4}\)

P (x = 4) = \(^{6}C_4 \times (p)^4 \times(1-p)^{6-4}\)

P (x = 4) = \({{6!}\over {{(6-4)!}4!}} \times (p)^4 \times(1-p)^{2}\)

P (x = 4) = \(15 \times (p)^4 \times(1-p)^{2}\)

As, 9P(X = 4) = P(X = 2)

\(9\times 15 \times (p)^4 \times(1-p)^{2} = ​​15 \times (p)^2 \times(1-p)^{4}\)

\(9\times (p)^2 = (1-p)^{2}\)

\(9 = ({{1-p}\over p})^2\)

\(3 = {{1-p}\over p}\)

3p = 1 - p

p = \( {{1}\over 4}\)

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 596

National Defence Academy (NDA) 2023• Mathematics• Mathematics

Consider the following for the next items that follow:

Three boys P, Q, R and three girls S, T, U are to be arranged in a row for a group photograph.

What is the probability that all three boys sit together?

  1. AOption (a)

    \(\frac{1}{5}\)

  2. BOption (b)

    \(\frac{1}{4}\)

  3. COption (c)

    \(\frac{1}{3}\)

  4. DOption (d)

    \(\frac{1}{12}\)

View answer & explanation

Correct answer

Answer: (a)A.

\(\frac{1}{5}\)

Why this answer is correct

Given: Three boys P, Q, R and three girls S, T, U are to be arranged in a row for a group photograph.

Calculation:

The probability of all three boys sitting together:

Treat the boys as one unit (B).

Now, we have four units (B, S, T, U) to arrange. This can be done in 4! ways.

Within the boys' block (B), the three boys can arrange themselves in 3! ways.

Therefore, the total number of favorable outcomes in this case is 4! × 3!.

6 people can be arranged in a row in 6! ways.

\(Probability = {Favorable: outcomes \over Total :possible :outcomes}\)

Probability \(= {{(4! × 3!)} \over 6!}\)

Probability \(= {{(24 × 6)} \over 720}\)

Probability = 1/5

Therefore, the probability that all three boys sit together is 1/5.

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 597

National Defence Academy (NDA) 2023• Mathematics• Mathematics

Consider the following for the next items that follow:

Three boys P, Q, R and three girls S, T, U are to be arranged in a row for a group photograph.

What is the probability that boys and girls sit alternatively?

  1. AOption (a)

    \(\frac{4}{5}\)

  2. BOption (b)

    \(\frac{1}{10}\)

  3. COption (c)

    \(\frac{5}{6}\)

  4. DOption (d)

    \(\frac{1}{7}\)

View answer & explanation

Correct answer

Answer: (b)B.

\(\frac{1}{10}\)

Why this answer is correct

Given: Three boys P, Q, R and three girls S, T, U are to be arranged in a row for a group photograph.

Case 1: Starting with a Girl

In this case, boys can be arranged in this way: B_B_B

Boys can be arranged in this manner in 3! ways.

Girls can be arranged in this manner in 3! ways.

Total Favorable outcomes = 3! × 3! = 6 × 6 = 36 ways

Case 2: Starting with a Boy

In this case, boys can be arranged in this way: B_B_B_

Boys can be arranged in this manner in 3! ways.

Girls can be arranged in this manner in 3! ways.

Total Favorable outcomes = 3! × 3! = 6 × 6 = 36 ways

Total favorable outcomes (Case 1 + Case 2) = 36 + 36 = 72.

Total possible arrangements for 6 people: 6! = 720

The probability that boys and girls sit alternatively,

\(Probability = {Favorable: outcomes \over Total :possible :outcomes}\)

Probability \(= {{(72)} \over 720}\)

Probability = 1/10

Therefore, the probability that boys and girls sit alternatively is 1/10.

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 598

National Defence Academy (NDA) 2023• Mathematics• Mathematics

Consider the following for the next items that follow:

Three boys P, Q, R and three girls S, T, U are to be arranged in a row for a group photograph.

What is the probability that no two girls sit together?

  1. AOption (a)

    \(\frac{2}{5}\)

  2. BOption (b)

    \(\frac{3}{5}\)

  3. COption (c)

    \(\frac{1}{18}\)

  4. DOption (d)

    \(\frac{1}{5}\)

View answer & explanation

Correct answer

Answer: (d)D.

\(\frac{1}{5}\)

Why this answer is correct

Given: Three boys P, Q, R and three girls S, T, U are to be arranged in a row for a group photograph.

Calculation:

Boys can be arranged in this way: B_B_B

Boys can be arranged in this manner in 3! ways.

Girls can be arranged in this manner in 3! × 4P3 ways.

Total Favorable outcomes = 3! × 3! × 4P3 = 6 × 6 × 4! = 144 ways

The probability that no two girls sit together,

6 people can be arranged in a row in 6! i.e. 720 ways.

\(Probability = {Favorable: outcomes \over Total :possible :outcomes}\)

Probability \(= {{(144)} \over 720}\)

Probability = 1/5

Therefore, the probability that no two girls sit together is 1/5.

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 599

National Defence Academy (NDA) 2023• Mathematics• Mathematics

Consider the following for the next items that follow:

Three boys P, Q, R and three girls S, T, U are to be arranged in a row for a group photograph.

What is the probability that P and Q take the two end positions?

  1. AOption (a)

    \(\frac{1}{15}\)

  2. BOption (b)

    \(\frac{7}{15}\)

  3. COption (c)

    \(\frac{14}{15}\)

  4. DOption (d)

    \(\frac{11}{45}\)

View answer & explanation

Correct answer

Answer: (a)A.

\(\frac{1}{15}\)

Why this answer is correct

Given: Three boys P, Q, R and three girls S, T, U are to be arranged in a row for a group photograph.

Calculation:

P and Q take the two end positions: P_ _ _ _ Q or Q_ _ _ _ P

4 people arrange between P and Q in 4! ways.

Total Favorable outcomes = 2 × 4! = 2 × 24 = 48 ways

The probability that P and Q take the two end positions,

6 people can be arranged in a row in 6! ways.

\(Probability = {Favorable: outcomes \over Total :possible :outcomes}\)

Probability \(= {{(48)} \over 720}\)

Probability = 1/15

Therefore, the probability that P and Q take the two end positions is 1/15.

NDA Mathematics 3 Sep 2023 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 600

National Defence Academy (NDA) 2023• Mathematics• Mathematics

Consider the following for the next items that follow:

Three boys P, Q, R and three girls S, T, U are to be arranged in a row for a group photograph.

What is the probability that Q and U sit together?

  1. AOption (a)

    \(\frac{2}{3}\)

  2. BOption (b)

    \(\frac{1}{4}\)

  3. COption (c)

    \(\frac{5}{6}\)

  4. DOption (d)

    \(\frac{1}{3}\)

View answer & explanation

Correct answer

Answer: (d)D.

\(\frac{1}{3}\)

Why this answer is correct

Given: Three boys P, Q, R and three girls S, T, U are to be arranged in a row for a group photograph.

Calculation:

The probability that Q and U sit together.

Treat the Q and U as one unit (X).

Now, we have five units (X, P, R, S, T) to arrange. This can be done in 5! ways.

Within Q and U, they can arrange themselves in 2! ways.

Therefore, the total number of favorable outcomes in this case is 5! × 2!.

6 people can be arranged in a row in 6! ways.

\(Probability = {Favorable: outcomes \over Total :possible :outcomes}\)

Probability \(= {{(5! × 2!)} \over 6!}\)

Probability \(= {{(120 × 2)} \over 720}\)

Probability = 1/3

Therefore, the probability that all three boys sit together is 1/3.

How should I prepare for Quantitative Aptitude in National Defence Academy (NDA)?

Solving previous year questions by topic helps identify recurring concepts, question patterns, and frequently asked sub-topics. Attempt each question, review the step-by-step explanation, and note down key formulas or facts for quick revision before the exam.